OpenStudy (anonymous):

What is the sum of the geometric series? 15,658 6,138 12,276 756

OpenStudy (anonymous):



OpenStudy (anonymous):

@jhonyy9 @abb0t @sourwing @Destinymasha @whpalmer4 ?? Help :) Anyone...

OpenStudy (abb0t):

Use a calculator. Lol. You're basicalling finding the sum of 6[2\(^1\)+2\(^2\)+2\(^3\)+...+2\(^{10}\)]

OpenStudy (anonymous):

Oh so I just add that all up till 2 to the power of 10?

OpenStudy (abb0t):

Well, from 2\(^1\) until you get to 2\(^{10}\) and then multiplying it by 6.

OpenStudy (anonymous):

No wonder I keep getting wrong answer... Can u wait till I solve, so u can double check my answer :)

OpenStudy (abb0t):

I don't have a calculator with me.

OpenStudy (anonymous):

I added them all together which is (2+4+8+16+32+64+128+256+512+1024) = 2046 * 6 = 12,276 Is this correct ??

OpenStudy (abb0t):

It looks like it matches one of the given anser choices, so probably.

OpenStudy (anonymous):

@abb0t does this look correct? :) Thanks for helping :)

OpenStudy (abb0t):

Yes. Your process looks correct.

OpenStudy (anonymous):

Thanks!!! :D